OpenStudy (anonymous):

NEED FAST ANSWER AND WILL FAN AND GIVE MEDAL Cell Phone Company A charges $20 each month plus $0.03 per text. Cell Phone Company B charges $5 each month plus $0.07 per text. Write a system of equations to model the situation using c for cost and t for number of texts. How many texts does a person need to send in a month to make the costs from both companies equal?

OpenStudy (triciaal):

company A: C= 20 +0.03t company B: C = 5 + 0.07t set the cost equal and solve for t

OpenStudy (triciaal):

t = 375 to check answer 5 + 0.07(375) = Cost for B 20 + 0.03(375) = Cost for A cost for B = Cost for A = 31.25 so number of texts = 375